OpenStudy (anonymous):

Hi simple and easy Math question ! Medal and fan awarded :) Below are the steps to solve an equation: Step 1: |x - 2| + 3 = 7 Step 2: |x - 2| = 7 - 3 Step 3: |x - 2| = 4 Which of the following is a correct next step to solve the equation? *answers below*

OpenStudy (anonymous):

x + 2 = -4 -x – 2 = 4 x + 2 = 4 x - 2 = -4

OpenStudy (anonymous):

The next two correct steps would be x - 2 = 4 (positive) and x - 2 = -4 (negative) so the fourth choice is the correct one.

OpenStudy (anonymous):

Okay cool thanks